### v1

@@ Abstract
The value of second-order methods lies in the use of curvature information. Yet, this information is costly to extract and once obtained, valuable negative curvature information is often discarded so that the method is globally convergent. This limits the effectiveness of second-order methods in modern machine learning. In this paper, we show that second-order and second-order-like methods are promising optimizers for neural networks provided that we add one ingredient: negative step sizes. We show that under very general conditions, methods that produce ascent directions are globally convergent when combined with a Wolfe line search that allows both positive and negative step sizes. We experimentally demonstrate that using negative step sizes is often more effective than common Hessian modification methods.
@@ 1 Introduction
Modern machine learning typically involves training deep neural networks and is thus a high-dimensional and non-convex problem. Non-convexity makes learning difficult for second-order methods that are attracted to saddle points and local maxima. By and large, gradient based methods such as gradient descent with momentum, Adam (Kingma and Ba, 2014) and RMSprop (Tieleman and Hinton, 2012) are seen as the only viable optimizers in machine learning. This is a source of frustration. Newton’s method and second-order-like methods, such as limited-memory quasi-Newton (QN) methods, excel in traditional machine learning but can fail to even converge in the deep learning setting. First-order methods converge but slowly, as good convergence progress can only be expected if curvature information of the loss landscape is used. We want optimizers that use negative curvature information but these methods may yield search directions that point the wrong way.
Many variants of second-order methods exist that ensure global convergence by enforcing a positive-definite second-order approximation. This approach may not be accurate for non-convex problems where the Hessian may have negative eigenvalues. Other methods, such as the dogleg method (Powell, 1970) and trust-region methods, combine gradient descent (GD) with a second-order method to ensure progress. Yet, for problems that are ill-conditioned and non-convex, GD ensures progress but at an extremely slow rate (Martens, 2010) and the second-order direction could point uphill or to saddle points. In such cases, simply taking the negative of the second-order direction may be a better choice, as this points downhill while maintaining all second-order information (see Figure 1).
To the best of our knowledge, using negative step sizes is relatively unexplored in optimization. We found one mention of negative step sizes in an analysis of the Fletcher-Reeves Fletcher and Reeves (1964) non-linear conjugate gradient method Dai (1999) for differentiable and Lipschitz-smooth objective functions.
@@ 1.1 Contribution
This paper examines the largely unexplored role of negative step sizes in optimization and show that taking a backward step is a computationally inexpensive way to incorporate negative curvature information. Our experiments suggest that a quasi-Newton (QN) method, symmetric rank 1 (SR1) Conn et al. (1991), combined with negative step sizes, may be an overlooked method useful for training neural networks. This is achieved without using common approaches to “fix” second-order methods [^1] In this paper, we refer to second-order-like methods that approximate the Hessian, such as QN methods, as “second-order methods”. This is technically incorrect because these methods do not compute second derivatives. This misuse of terminology is for readability. such as Hessian modifications. We also show that extending the Wolfe line search to both positive and negative step sizes ensures that second-order methods satisfy the Zoutendijk condition. Thus, second-order methods combined with good step size choices are globally convergent even in non-convex settings.
Figure 1: In this two-dimensional non-convex minimization problem, the negative of the Newton’s direction (green) is a better search direction than the gradient descent direction (blue), Newton’s direction (red) or any convex combination of both.
@@ 2 Common approaches for globally convergent second-order methods
There are many variants of second-order method that modify the (approximate) Hessian for positive definiteness and thus ensure a descent direction (see Sections 3.4 of Nocedal and Wright (2006)). A spectral decomposition will find the negative eigenvalues. One could then zero out the negative eigenvalues, flip the sign of the negative eigenvalues Paternain et al. (2019) or shift the Hessian by subtracting a diagonal matrix that contains the most negative eigenvalue of the Hessian (also known as damping). These modifications, however, require an expensive eigenvalue decomposition and may no longer reflect accurate curvature information. Removing negative eigenvalues could also be detrimental for high-dimensional non-convex problems because following paths of negative curvature is a way to escape saddle points Choromanska et al. (2015); Moré and Sorensen (1979).
@@
Trust-region approaches
Trust region (TR) methods is an option for obtaining global convergence and, unlike line searches, they work with methods that produce ascent directions Conn et al. (2000). Newton’s and QN methods often employ a TR approach that minimizes the model function over a two-dimensional subspace Sorensen (1982); Byrd et al. (1988); Schultz et al. (1985); Ramanmurthy and Duffy (2016). Saddle-free Newton’s method Pascanu et al. (2014); Dauphin et al. (2014), Newton’s method with generalized TR, was proposed as a version of Newton’s method that can escape saddle points. However, TR is comparable to damping Moré and Sorensen (1983) and thus loses curvature information in the same way as Hessian modifications.
Cubic regularization
Cubic-regularization (CR) uses an additional third-order term so that the second-order method is not attracted to local maxima and saddle points. Originally introduced for Newton’s method Nesterov and Polyak (2006), CR was recently extended to QN methods (Kamzolov et al., 2023; Benson and Shanno, 2018). However, CR generally requires special solvers for its third-order subproblem.
Positive definite-enforcing updates
Some QN methods, such as BFGS (Broyden, 1967; Fletcher and Reeves, 1964; Goldfarb, 1970; Shanno, 1970), are designed to guarantee positive definiteness of the Hessian approximation at every update step as long as the step size used satisfies the Wolfe conditions (Wolfe, 1969). It has been suggested, however, that QN variants that do not enforce positive definiteness, such as SR1, converge faster by adhering closer to the true Hessian (Conn et al., 1991; Khalfan et al., 1993). Other updates that enforce positive definiteness include natural gradient methods Amari (1998) that use the covariance matrix of the gradients instead of the Hessian, Gauss-Newton methods and Kronecker-factored approximate curvature (KFAC) Martens and Grosse (2015). These methods, however, may also sacrifice accurate curvature information by enforcing positive definiteness.
3 Global convergence of ascent-descent methods
Even though negative step sizes may result in a larger decrease on some iterations, we may be concerned that the algorithm might not converge. In this section, we give a general result showing that methods that may produce ascent directions and that allow negative step sizes converge under very general conditions. Our goal is to minimize a function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ without constraints. $f$ is assumed to be twice differentiable, Lipschitz-smooth and bounded below. We use a deterministic iterative algorithm that, on the $k$th iteration, calculates a search direction ${\bm{p}}_{k}$ and step size $\alpha_{k}$, and updates its iterate ${\bm{x}}$ by
${\bm{x}}_{k+1}={\bm{x}}_{k}+\alpha_{k}{\bm{p}}_{k}.$ (1)
@@ 3 Global convergence of ascent-descent methods
Proposition 3.1.
Consider any algorithm with iteration of the form equation 1 where ${\bm{p}}_{k}$ is direction that is not orthogonal to the gradient direction, and non-zero step sizes $|\alpha_{k}|\geq\epsilon>0$ satisfy the Wolfe conditions (6 and 7). Suppose that $f$ is bounded below in $\mathcal{R}^{n}$ and that $f$ is continuously differentiable in an open set $\mathcal{N}$ containing the level set $\triangleq$$\mathcal{L}\triangleq\{{\bm{x}}:f({\bm{x}})\leq f({\bm{x}}_{0})\}$, where ${\bm{x}}_{0}$ is the initial iterate. Assume also that the gradient $\nabla f$ is Lipschitz continuous on $\mathcal{N}$. Then, the algorithm satisfies Zoutendijk’s condition (4) is therefore globally convergent (2).
4 Experiments
4.1 QN methods and neural network training
Because the objective is non-convex, the true Hessian at iterate ${\bm{x}}_{k}$ may not be positive definite. Thus, SR1 produces Hessian approximations that are not positive definite and a ${\bm{p}}_{k}$ that does not point downhill. With a regular line search, SR1 diverges within the first few steps (blue). With negative step sizes, SR1 becomes the an effective optimizer. This effect appears to be more pronounced as the number of network layers increase. BFGS combined with a Wolfe line search guarantees positive definiteness in its Hessian update and thus never diverges even with positive only step sizes. Yet, a positive definite Hessian may not accurately capture the optimization landscape. The lack of negative curvature information in the BFGS Hessian approximation may be the reason its training error is higher than that of SR1 after the maximum number of iterations is reached.
For further details, please refer to Appendix D.
4.2 Limited-memory QN methods and neural network training
Table 1 shows the result of using two limited-memory QN methods, l-BFGS and l-SR1, to train a neural network across different datasets. Gradient descent (GD) and Adam are plotted for comparison. The purpose of this experiment is to compare two globalization strategies: damping and negative step sizes. Damping requires an eigenvalue decomposition and is therefore an expensive operation. Thus, in many cases where the dataset has a large number of features and the Hessian is large (e.g. colon-cancer), l-SR1 combined with damping did not finish in the allocated amount of time (shown as DNF). l-BFGS combined with a Wolfe line search does not require damping because its Hessian approximations are guaranteed to be positive definite. Yet, a lack of negative curvature information may hurt its effectiveness in training neural networks, and it tends to perform worse than l-SR1 with negative step sizes.

### v2

@@ Abstract
The value of second-order methods lies in the use of curvature information. Yet, this information is costly to extract and once obtained, standard methods to achieve global convergence often discard valuable negative curvature information. This limits the effectiveness of second-order methods in modern machine learning. In this paper, we show that second-order and second-order-like methods are promising optimizers for neural networks provided that we add one ingredient: negative step sizes. We show that under very general conditions, methods that produce ascent directions are globally convergent when combined with a Wolfe line search that allows both positive and negative step sizes. We experimentally demonstrate that using negative step sizes is often more effective than common Hessian modification methods.
@@ 1 Introduction
Training neural networks involves minimizing a non-convex objective. Often, second-order methods are considered unsuitable for this task because they are attracted to saddle points and local maxima. By and large, gradient based methods such as gradient descent (GD) with momentum, Adam (Kingma and Ba, 2014) and RMSprop (Tieleman and Hinton, 2012) are seen as the only viable optimizers in machine learning. This is a source of frustration. Newton’s method and second-order-like methods, such as limited-memory quasi-Newton (QN) methods, excel in traditional machine learning but can fail to even converge in the deep learning setting. First-order methods converge but slowly, as good convergence progress can only be expected if curvature information of the loss landscape is used. We want optimizers that use negative curvature information but these methods may yield search directions that point the wrong way.
Many variants of second-order methods ensure global convergence. Popular approaches include Hessian modifications, trust-region methods and cubic regularization (see Section 2 for a short summary). These methods use the same trick: align the search direction closer to the gradient direction to ensure descent. Yet, this approach could slow down optimization progress. For problems that are ill-conditioned and non-convex, GD ensures progress but at an extremely slow rate (Martens, 2010). When the second-order direction points uphill, simply taking the opposite direction may be a better choice than moving the step closer to the gradient step. The negative direction points downhill and maintains all second-order information (see Figure 1). To the best of our knowledge, using negative step sizes is relatively unexplored in optimization. We found one mention of negative step sizes in an analysis of the Fletcher-Reeves Fletcher and Reeves (1964) non-linear conjugate gradient method Dai (1999) for differentiable and Lipschitz-smooth objective functions.
@@ 1.1 Contribution
This paper examines the largely unexplored role of negative step sizes in optimization and shows that taking a backward step is a computationally inexpensive way to incorporate negative curvature information. Our experiments suggest that a quasi-Newton (QN) method, symmetric rank one (SR1) Conn et al. (1991), combined with negative step sizes, may be an overlooked method useful for training neural networks. This is achieved without using common approaches to “fix” second-order methods [^1] In this paper, we refer to second-order-like methods that approximate the Hessian, such as QN methods, as “second-order methods”. This is technically incorrect because these methods do not compute second derivatives. This misuse of terminology is for readability. such as Hessian modifications. We also show that extending the Wolfe line search to both positive and negative step sizes ensures that second-order methods satisfy the Zoutendijk condition. Thus, second-order methods combined with good step size choices are globally convergent even in non-convex settings.
Figure 1: In this two-dimensional non-convex minimization problem, the negative of the Newton’s direction (green) is a better search direction than the gradient descent direction (blue), Newton’s direction (red) or any non-negative combination of both.
@@ 2 Common approaches for globally convergent second-order methods
There are many variants of second-order method that modify the (approximate) Hessian for positive definiteness and thus ensure a descent direction (see Sections 3.4 of Nocedal and Wright (2006)). A spectral decomposition will find the negative eigenvalues. One could then replace the negative eigenvalues by small positive values, flip the sign of the negative eigenvalues Paternain et al. (2019) or shift the Hessian by subtracting a diagonal matrix that contains the most negative eigenvalue of the Hessian (also known as damping). These modifications, however, require an expensive eigenvalue decomposition and may no longer reflect accurate curvature information. Removing negative eigenvalues could also be detrimental for high-dimensional non-convex problems because following paths of negative curvature is a way to escape saddle points Choromanska et al. (2015); Moré and Sorensen (1979).
Trust-region approaches
Trust region (TR) methods are an option for obtaining global convergence and, unlike line searches, they work with methods that produce ascent directions Conn et al. (2000). Variants of TR methods include the Leveberg-Marquadt Levenberg (1944); Marquardt (1963) and dogleg (Powell, 1970) methods. Newton’s and QN methods often employ a TR approach that minimizes the model function over a two-dimensional subspace Sorensen (1982); Byrd et al. (1988); Schultz et al. (1985); Ramanmurthy and Duffy (2016). Saddle-free Newton’s method Pascanu et al. (2014); Dauphin et al. (2014), Newton’s method with generalized TR, was proposed as a version of Newton’s method that can escape saddle points. However, because the solution to the TR subproblem is equivalent to adding a positive definite diagonal matrix to the (approximate) Hessian Moré and Sorensen (1983), this approach also loses curvature information. TR conjugate gradient methods Steihaug (1983) use directions of negative curvature, but this also combines the original search direction with the negative gradient direction.
Cubic regularization
Cubic-regularization (CR) uses an additional third-order term so that the second-order method is not attracted to local maxima and saddle points. Originally introduced for Newton’s method Nesterov and Polyak (2006), CR was recently extended to QN methods (Kamzolov et al., 2023; Benson and Shanno, 2018). However, CR generally requires special solvers for its third-order subproblem. CR is equivalent to using a Hessian that is modified to be positive definite Malitsky and Mishchenko (2020).
Positive definite-enforcing updates
Some QN methods, such as BFGS (Broyden, 1967; Fletcher and Reeves, 1964; Goldfarb, 1970; Shanno, 1970), are designed to guarantee positive definiteness of the Hessian approximation at every update step as long as the step size used satisfies the Wolfe conditions (Wolfe, 1969). It has been suggested, however, that QN variants that do not enforce positive definiteness, such as SR1, converge faster by adhering closer to the true Hessian (Conn et al., 1991; Khalfan et al., 1993). Other updates that enforce positive (semi-) definiteness include natural gradient methods Amari (1998) that use the covariance matrix of the gradients instead of the Hessian, Gauss-Newton methods and Kronecker-factored approximate curvature (KFAC) Martens and Grosse (2015). These methods, however, may also sacrifice accurate curvature information by enforcing positive (semi-) definiteness.
3 Global convergence of ascent-descent methods
Even though negative step sizes may result in a larger decrease on some iterations, we may be concerned that the algorithm might not converge. In this section, we give a general result showing that methods that may produce ascent directions and that allow negative step sizes converge under very general conditions. Our goal is to minimize a function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ without constraints. The function $f$ is assumed to be twice differentiable, Lipschitz-smooth and bounded below. We use a deterministic iterative algorithm that, on the $k$th iteration, calculates a search direction ${\bm{p}}_{k}$ and step size $\alpha_{k}$, and updates its iterate ${\bm{x}}$ by
${\bm{x}}_{k+1}={\bm{x}}_{k}+\alpha_{k}{\bm{p}}_{k}.$ (1)
Proposition 3.1.
Consider any algorithm with updates of the form equation 1 where ${\bm{p}}_{k}$ is a direction that is not orthogonal to the gradient direction, and non-zero step sizes $|\alpha_{k}|\geq\epsilon>0$ satisfy the Wolfe conditions (6 and 7). Suppose that $f$ is bounded below in $\mathcal{R}^{n}$ and that $f$ is continuously differentiable in an open set $\mathcal{N}$ containing the level set $\triangleq$$\mathcal{L}\triangleq\{{\bm{x}}:f({\bm{x}})\leq f({\bm{x}}_{0})\}$, where ${\bm{x}}_{0}$ is the initial iterate. Assume also that the gradient $\nabla f$ is Lipschitz continuous on $\mathcal{N}$. Then, the algorithm satisfies Zoutendijk’s condition (4) is therefore globally convergent (2).
4 Experiments
@@
4.1 QN methods and neural network training
Because the objective is non-convex, the true Hessian at iterate ${\bm{x}}_{k}$ may not be positive definite. Thus, SR1 produces Hessian approximations that are not positive definite and a ${\bm{p}}_{k}$ that does not point downhill. With a regular line search, SR1 diverges within the first few steps (blue). With negative step sizes, SR1 becomes the an effective optimizer. This effect appears to be more pronounced as the number of network layers increase. BFGS combined with a Wolfe line search guarantees positive definiteness in its Hessian update and thus never diverges even with positive only step sizes. Yet, a positive definite Hessian may not accurately capture the optimization landscape. The lack of negative curvature information in the BFGS Hessian approximation may be the reason its training error is higher than that of SR1 after the maximum number of iterations is reached. For further details, please refer to Appendix D.
4.2 Limited-memory QN methods and neural network training
@@
Table 1 shows the result of using two limited-memory QN methods, l-BFGS Liu and Nocedal (1989) and l-SR1 Bryd et al. (1994), to train a neural network across different datasets. Gradient descent (GD) and Adam are plotted for comparison. The purpose of this experiment is to compare two globalization strategies: damping and negative step sizes. Damping requires an eigenvalue decomposition and is therefore an expensive operation. Thus, in many cases where the dataset has a large number of features and the Hessian is large (e.g. colon-cancer), l-SR1 combined with damping did not finish in the allocated amount of time (shown as DNF). l-BFGS combined with a Wolfe line search does not require damping because its Hessian approximations are guaranteed to be positive definite. Yet, a lack of negative curvature information may hurt its effectiveness in training neural networks, and it tends to perform worse than l-SR1 with negative step sizes.